OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7.

OpenStudy (anonymous):

since it's parallel to y=x+7, it must have the same slope of 1. Now you have a slope (m=1) , and a point (6, -2) , you can put these into the point-slope form for the equation of a line: y - y1 = m(x - x1)